whats is the minimum or maximum of y=x^2+8x+17
Answer by lwsshak3(11628)

whats is the minimum or maximum of y=x^2+8x+17
Given equation is that of a parabola that opens upwards, so it has a minimum.
complete the square:
y=(x^2+8x+16)-16+17
y=(x+4)^2+1
minimum=1 when x=-4